Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

Leinster Indoor Prone Rifle Open

Distribution

Dr. Trevor West DUCAC Chairman
Aubrey Storey DUCAC Secretary
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Tony O’Donnell Student Record
Susan Thompson Trinity News

Summary

Venue Rathdrum Rifle and Pistol Club
Date November 4th, 2001
Discipline 25yd Prone Rifle

DURC Attendees

Mark Dennehy  John Keeney
Peter Byrne  Deirdre Ninaber
Adrian Costigan  Eric Farrell
Michael Madigan  Shane Kelly

Alan Smyth

Scores

John Keeney 581
Alan Smyth 578
Deirdre Ninaber 568
Peter Byrne 564
Adrian Costigan 550
Eric Farrell 550
Mark Dennehy 549
Shane Kelly 535
Michael Madigan 518

Prizes Won

John Keeney Second place overall

First place, Class B

Alan Smyth First place, Class C
Deirdre Ninaber First place, Class D

Report

DURC sent a squad to the last open smallbore competition of the calendar year to Rathdrum for the Leinster Indoor Open. The early start took it’s toll on some - Mark Dennehy, the current National Champion in 25yd prone rifle, shot the worst score he has shoot indoors since he began competing eight years ago. However, he felt that this was justified by having spent the previous two days representing the National Governing Body of target shooting (the National Rifle and Pistol Association of Ireland) at the Sports Development Conference, where the Irish Sports Council and the Sports Council for Northern Ireland set out the path they wanted Sport in Ireland to progress along for the next decade.

John Keeney, however, proved far more resiliant and despite having travelled from Kildare only the night before, shot a 581 to take second place overall, beaten only by Dave Cullen, a 40-year veteran of target shooting, by a mere 5 points and on Dave’s home range too! John also took the first place prize in class B.

In face, DURC swept the boards in the classification prizes - first place in Classes B, C and D went to DURC shooters John Keeney, Alan Smyth and Deirdre Ninaber. Deirdre was in fact competing in only her second competition ever - having only ever shot once before (in the 25yd Nationals) scoring a 491. Therefore in one competition she has improved by 77 points - an incredible leap in performance which hasn’t been seen in DURC in many years. We expect big things of Deirdre in next year’s Target Rifle Season, which begins in March, and possibly in the annual invitational shoot in RRPC at christmas as well!

John Keeney and Deirdre Ninaber accepting their prizes

Misé le meas,

Mark Dennehy, DURC Secretary